# Case Law Grounding: Using Precedents to Align Decision-Making for Humans and AI

### v3

@@ Abstract.
Communities and groups often need to make decisions grounded by social norms and preferences, such as when moderating content or providing judgments for aligning AI systems. Prevailing approaches to provide this grounding have primarily centered around constructing high-level guidelines and criteria, similar to legal “constitutions”. However, it can be challenging to specify social norms and preferences consistently and accurately through constitutions alone. In this work, we take inspiration from legal systems and introduce “case law grounding” (CLG)—a novel approach for grounding decision-making that uses past cases and decisions (precedents) to ground future decisions in a way that can be utilized by human-led processes or implemented through prompting large language models (LLMs). We evaluate how accurately CLG grounds decisions with five groups and communities spread across two decision task domains, comparing against a traditional constitutional grounding approach, and find that in 4 out of 5 groups, decisions produced with CLG were significantly more accurately aligned to ground truth: 16.0–23.3 %-points higher accuracy using the human-led process, and 20.8–32.9 %-points higher when prompting LLMs. We also evaluate the impact of different configurations of CLG, such as the case retrieval window size and whether to enforce binding decisions based on selected precedents, showing support for using binding decisions and preferring larger retrieval windows. Finally, we discuss the limitations of our case-based approach as well as how it may be best used to augment existing constitutional approaches when it comes to aligning human and AI decisions.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Many technological systems today are entangled in social decision-making situations. Online spaces for inter-personal communication, like social media platforms, need to contend with decisions around how to moderate content in a way that maintains a healthy community (Gongane et al., 2022). General purpose AI assistants, like LLM-backed chatbots, also need to contend with decisions around how to generate answers in a way that users find appropriate (Ji et al., 2023; Shen et al., 2023; Gabriel, 2020). The scale of these systems means that getting the decisions wrong can have major negative consequences for people (Haimson et al., 2021; Weidinger et al., 2021).
Unlike systems that work with more objective determinations, decisions around social concepts are especially hard because: (1) between groups or communities, there often isn’t a shared consensus around what is “correct”, and (2) within a group or community, norms and expectations are often implied and not formalized. This makes it hard to make consistently good decisions at scale—groups and communities either make efforts to ensure decision quality through highly involved procedures (e.g., deliberation (Im et al., 2018; Small et al., 2021), digital juries (Fan and Zhang, 2020), expert panels (Klonick, 2020)), or they might trade off some decision quality for scale (e.g., building classifiers and auto-moderation (Jhaver et al., 2019b; Wulczyn et al., 2017), outsourcing to paid crowd workers (Satariano and Isaac, 2021)).
@@
One way that current systems have attempted to bridge this demand has been to try and distill social norms and preferences into general decision criteria. For example, moderators of communities on social media platforms often create community guidelines that establish what content is allowable, and how violations are to be dealt with (Chandrasekharan et al., 2018a; Chandrasekharan and Gilbert, 2019). Similarly, builders and operators of AI assistant services often come up with principles (sometimes with help from the general public (Huang et al., 2024)), that are then used to synthesize training and fine-tuning examples (Bai et al., 2022) or become part of the prompts (Lin et al., 2024) for models. However, in practice, imprecise language can leave room for ambiguity (Fiesler et al., 2018; Chen et al., 2021) around edge cases, and varied decision-makers can result in divergent interpretations (Gordon et al., 2021). Is there another way we could better support making socially-aligned decisions at scale? For a solution, we look to the realm of legal systems: In legal systems, the creation of principles and guidelines to ground decisions parallels that of creating and applying constitutional law. However, many legal systems also utilize a different way to ground more nuanced decisions—case law. With case law, rather than relying on lawmakers (or, in our case, moderators and experts) to resolve every edge case and ambiguity through modifications and additions to the constitution, decision-makers can also rely on past decisions as sources of clarification and disambiguation.
@@
Taking inspiration from case law legal systems, we introduce the case law grounding (CLG), a general workflow for grounding decisions on subjective or socially-aligned tasks that can be support human-led or LLM prompting processes. Like with case law in the legal realm, CLG focuses on making use of precedents (past decisions) as the primary source of guidance for new decisions. CLG accomplishes this through 3 steps: (1) case retrieval, where past cases potentially relevant to the current decision are retrieved; (2) precedent selection, where retrieved precedents are further refined for applicability to the current decision; and finally (3) decision synthesis, where the selected precedents and decisions are aggregated and synthesized into a decision for the new case. As online groups and communities often need to make these types of decisions at-scale, we also examine both ways in which the CLG workflow could be implemented: the human-led version targeting use cases where members within the group or community are the decision-makers; and an LLM-prompting based version targeting use cases where groups and communities may want to scale up with the help of AI while also ensuring decisions remain aligned to their own values.
@@
Recruiting participants from 3 Reddit communities and 2 demographically-recruited population groups, we evaluated both human against a traditional workflow that uses constitutional grounding to support decisions. In the human-led process, participants made binary “keep”/“remove” decisions around content moderation outcomes (Park et al., 2021) within a community they frequent, or rated toxicity of social media posts (Kumar et al., 2021) for their demographic group, recruiting $n=35$ Redditors and $n=60$ crowd workers for the mod and tox domains respectively. We also evaluated an automated version of CLG for the same task domains, implemented through prompting an LLM.
In summary, in this work we make the following contributions:
  • We introduce a general framework for supporting group-aligned decision-making in subjective and social contexts, Case Law Grounding (CLG), and detail how CLG can be applied to two decision-making situations: a human-led process where community and group members make decisions grounded by precedents, and an automated process where an LLM is prompted to evaluate precedents instead of human annotators.
@@
  • We evaluate CLG with participants drawn from 5 groups (based on communities or demographic population groups) across two decision task domains, comparing against a more traditional approach to grounding decisions via constitutional rules and guidelines. We find that:
@@
    • RQ1: When using either the human-led or LLM prompting process, CLG was able to produce final decisions that more accurately matched observed ground truth compared to decisions grounded with constitutions, with statistically significant improvements observed in 4 out of 5 groups.
@@
    • RQ2a: Examining the retrieval process by simulating smaller retrieval window sizes, we find that for most domains, larger retrieval windows resulted in higher accuracy, however scaling characteristics varied across groups.
    • RQ2b: Comparing setups where precedents were binding versus those where they only served as examples, we find that the binding nature of precedents is important—in the 4 groups from RQ1 where the binding configuration produced statistically significant improvements in decision accuracy, the non-binding version did not produce statistically significant differences.
    • RQ3: We find that beyond the human-led process, LLMs were able to scale up to larger datasets, while still producing statistically significant improvements across all 6 tested groups.
  • Finally, we discuss limitations of human-led decision-making with case law grounding that arise from human evaluation of concrete cases and discuss how constitutional and case law grounding can complement each other in supporting group-aligned subjective decision-making.
@@ 2. Related Work
In this section, we discuss how our work relates to prior work around decision-making in social or subjective settings, constitutional approaches for consistent decision-making, and precedents of using cases to aid decision-making in prior social and technical systems.
2.1. Constitutions for Social and Subjective Concepts
The rise of online platforms has led to an increased demand for making socially-situated decisions through technological systems. For example, online communities such as Wikipedia, Reddit, and Facebook all need to make moderation decisions around what content is objectionable and should be removed (Van Aken et al., 2018; Kumar et al., 2021; Bhuiyan et al., 2020; Ross et al., 2017). However, the definitions of these concepts, like the criteria for moderation, can differ significantly between groups and communities (Chandrasekharan and Gilbert, 2019; Chandrasekharan et al., 2018a), which means it is important to ground decisions relevant to these groups locally to each group’s own preferences rather than universally across the entire platform.
2.1.1. Challenges of Creating Constitutions
Constitutions as Guidelines: Smaller online communities, such as subreddits, forums, or Discord servers, often rely on volunteer moderators drawn from members of the community to manually make moderation decisions (Matias, 2019). In these cases, moderators might already start with an intuitive understanding of what the criteria are simply based on their experience in the community. However, the socially-constructed intuition of what is desirable might not be aligned across all community members (Weld et al., 2021), and can lead to user frustration around decisions (Jhaver et al., 2019a). Community guidelines can codify these intuitive preferences as a constitution, which can then communicate criteria around moderation to all members.
However, in practice, moderators often have additional criteria undocumented in community guidelines, leading to mismatches when compared against real decisions (Juneja et al., 2020). It can also be difficult to neatly translate some concepts into actionable criteria than others, leading to guidelines that might still leave some concepts ambiguous and open to interpretation (Arhin et al., 2021; Fiesler et al., 2018). Finally, the high amount of effort that must be put into making clear and self-consistent guidelines (Im et al., 2018) leads to practical limits for many communities.
Constitutions as Rubrics: On bigger platforms that host content for millions or billions of people, consistency and transparency plays a more paramount role (Klonick, 2020). In these cases, organizations can easily justify effort to develop complex constitutions and even costs around training paid moderators to execute decisions based on the constitution (Roberts, 2014; Satariano and Isaac, 2021). One leaked set of internal manuals from Facebook contained over 1,400 pages across over 100 manuals (Guardian, 2017).
However, the extensive iteration often means that rubric-style constitutional grounding, while efficient, can be inflexible to context (Guardian, 2017; Press, 2024), which is especially undesirable given the social nature of such decisions. In cases where there are cultural or even sub-community level differences, it may be intractable to accommodate such nuances with rubric-style constitutional grounding (Jiang et al., 2021). When constitutions are relied on as the main way to guarantee consistent decisions, the resulting constitutions often become laundry lists of positive and negative carve-outs and examples for each category of content (Klonick, 2017).
Constitutions as Code and Models: Code and machine learning models can also be integrated into constitutional grounding when they are built with the intent of capturing some abstract aspect of the criteria. For example, classifiers can help identify and tag concepts (Wulczyn et al., 2017; Müller-Birn et al., 2013), which can be combined with constitutional guidelines to make automated decisions. Additionally, communities can build code-based constitutions to automatically execute decisions through systems like AutoModerator (Jhaver et al., 2019b), ORES (Halfaker and Geiger, 2019), and PolicyKit (Zhang et al., 2020).
However, building constitutions as code and models depends on accurate measurements of concepts that are often contextualized only in social settings and lack objective definitions. These types of annotations by nature come with annotator positionality (Chandrasekharan and Gilbert, 2019; Chandrasekharan et al., 2018a; Arhin et al., 2021; Denton et al., 2021) that result in conflicting or unclear decisions. The data being labeled (Chen et al., 2021) or annotation guidelines (Chang et al., 2017) can also be incomplete or ambiguous. Attempts to create objective instructions for annotation have also seen researchers and domain experts disagreeing on definitions (Fortuna et al., 2020) or criteria (Bhuiyan et al., 2020).
More recently, the advent of prompting-based control of large language models (LLMs) has also given rise to natural language constitutions being directly executed through system prompts in a zero-shot or few-shot way (Lin et al., 2024; Han, 2023). However, the limited control over how these models apply prompts means it can be difficult to build robust prompts that appropriately balance competing community desiderata (Raghavan, 2024).
2.1.2. Challenges of Using Constitutions for Decision-Making
Constitutions (in the form of rubrics and guidelines) have also been used extensively by crowdsourcing systems to improve annotation quality (Wu and Quinn, 2017), and thus many of the major improvements on making constitutional grounding more consistent can be found in systems for instructing the crowd.
One approach for improving consistency has been to continuously identify and refine problematic parts of rubrics associated with inconsistency, either automatically (Pradhan et al., 2022) or with worker cooperation (Bragg and Weld, 2018). Alternatively, other work has focused on real time training and evaluation during the decision-making phase of annotation, either in the form of gating (Liu et al., 2016) at the start, continuous evaluation during the task (Hettiachchi et al., 2021), or dynamic training (Werling et al., 2015).
However, applications of rubrics for consistency has been most beneficial in the area of more objective annotation tasks (Snow et al., 2008; Rashtchian et al., 2010). For primarily subjective surveys (Bentley et al., 2017) and annotations over socially-constructed concepts (Chen and Zhang, 2023), constructing comprehensive rubrics may be counterproductive (removing the subjectivity we would like to capture) or self-defeating (when we must come up with criteria that we are collecting judgments to define in the first place).
Because constitutional guidelines are created by humans, they can often omit aspects considered shared knowledge (Chandrasekharan et al., 2018b) in a community or seen as “common sense”, or fail to document criteria being applied in practice (Juneja et al., 2020). This can make it challenging for anyone without such context to accurately judge situations following guidelines (Jhaver et al., 2019a, c). Additionally, as norms and preferences shift over time (Lin et al., 2017), decisions based on past implicit knowledge may silently become out-of-date without easy means of correction. Together, this makes direct adaptation of constitutions to automated systems potentially fragile, leading to unexpected behaviors when the guidelines are under-defined or conflicting (Gordon et al., 2021).
Just like human decision-makers who may not always consistently and accurately interpret and decide based on constitutions, LLMs can fail in this respect as well, with prior work observing that LLMs grounded using constitutions can perform better than random for certain communities (Kumar et al., 2024).
2.2. Adopting Cases to Aid Judgments
The use of cases as judgment aids has been adopted by a wide variety of social and technological systems in the past. While sets of cases are not as compact or easily interpretable at a high level as constitutional policies, they can contribute to high-quality judgments by provide two types of support that complement deficiencies of constitutions: (1) grounding complex concepts—cases are specific and concrete and thus can illustrate concepts when abstract language fail; and (2) serving as a fine-grained policy definition—sets of cases with associated decisions can be adapted as a means to define fine-grained community- or group-specific policy via retrieval and interpolation.
2.2.1. Examples of Using Cases to Ground Complex Concepts
@@
Cases and case studies have broadly been used as a medium for reasoning around complex scenarios in various fields similar to the socially-constructed judgments we focus on. For example, moral case deliberation (Pierce, 2013; Sush and Najdowski, 2021; Emelin et al., 2021) is often used in medical settings to reflect on ethical situations (Molewijk et al., 2008), and case studies are common in business educational settings too (Farashahi and Tajeddin, 2018). In legal education, fictional cases are frequently used in exercises like moot court (Gaubatz, 1981; Ringel, 2004), as a medium for engaging in deliberative reasoning. Case-based reasoning (CBR) that focuses on analyzing the circumstances around specific case scenarios also has a long and rich history as a method for thinking about moral philosophy (Paulo, 2015) and in the realm of jurisprudence in the US legal system (Kagan, 2019).
Examples of input items and resulting judgments, often used to clarify rubrics and guidelines, can also be seen as a way of utilizing cases for grounding. In crowdsourced annotation, examples can augment text in rubrics to demonstrate or contrast how judgments should be made (Wu and Quinn, 2017), act as as tests to evaluate whether a concept or criteria was accurately understood (Liu et al., 2016; Hettiachchi et al., 2021; Singla et al., 2014), and even pose as reference anchors to ground the levels on a subjective scale (Chen et al., 2021).
@@
Finally, the detail provided by individual cases means that they can act as a medium for deliberation to reconcile differences and work towards shared understanding. We can see example of this in legal realm around adjudication (Carpenter, 1917), medical realm for diagnosis (Bertens et al., 2013; Schaekermann et al., 2019), in communities for working out moderation policy (Fan and Zhang, 2020; Pan et al., 2022), and more generally just around challenging tasks involving different lines of reasoning (Chen et al., 2019).
2.2.2. Case-based Reasoning and Data-centric Approaches to Defining Concepts
Case-based reasoning (CBR) has been applied in computational areas as well, forming a basis for computational reasoning systems in some classical AI systems (Aamodt and Plaza, 1994). Here, CBR is formalized as a 4-step process: (1) retrieve: where relevant “cases” (such as past problems and solutions) to a target problem are recalled; (2) reuse: where the reasoning and solutions from the retrieved cases are mapped to the new problem; (3) revise: where solutions are tested and any errors revised; and (4) retain: where the (revised) solution and problem are stored for use in future retrieval. However, application of CBR on more subjective tasks that require interpreting natural language has been limited due to limited capabilities of earlier models that focused on addressing objective reasoning tasks through formalized and symbolic approaches around what constitutes arguments and reasoning  (Ashley, 1989; Craw and Aamodt, 2018; Craw et al., 2001). In this work, we draw from general concepts in CBR around the process for making decisions with cases, while also focusing on adapting these ideas to subjective decision-making by (often non-expert) community members around socially-situated decisions based on norms and preferences which can be difficult (and possibly undesirable) to create symbolic formalizations of.
Viewed through a machine learning lens, cases also arise naturally as a data-centric view to define human criteria, including those that are subjective or dependent on social context (Simons et al., 2020; Jiang and de Marneffe, 2022). The training of classifiers using human annotated data is itself a form of grounding (a model’s) decisions through the medium of cases. However, viewing training data as an example of cases also shows the limitations of how we work with training data. For one, annotated datasets on socially-situated or subjective concepts are often produced by annotators who don’t have the correct context to judge, leading to biased results (Arhin et al., 2021; Denton et al., 2021). Additionally, aggregation used during pre-processing can hide disagreements that would have required human reconciliation (Gordon et al., 2021, 2022). Still, when data is relatively abundant, machine learning approaches tend to outperform expert constructed systems, though the best approaches depend on a combination of expert and learned aspects (Gennatas et al., 2020).
2.2.3. LLM Prompting as a Parallel of Human Judgment
Recent advances in natural language processing have brought us large language models (LLMs), which can act as general purpose problem solving systems that can be interacted with natural language prompts (Shi et al., 2023; Shen et al., 2023). The increasing adoption of LLMs in sociotechnical systems has presented both challenges and opportunities for decision-making at scale in social settings.
In recent works, LLMs have been adapted to complete a variety of complex reasoning and decision tasks (Minaee et al., 2024; Taori et al., 2023). Unlike traditional machine learning classification, where models produce decisions opaquely in a way that would differ significantly from a precedent reasoning process done by humans, the new capability of being able to work with natural language instructions and inputs makes it possible to create a close parallel of the human decision-making process through the use of LLM prompting. Along these lines, prompting strategies like Chain-of-Thought prompting (Wei et al., 2022) and prompt chains (Wu et al., 2021) have been proposed to improve the quality of model outputs that involve reasoning around a task. While other works have revisited ideas from crowdsourcing, breaking larger tasks into sub-tasks and applying quality control feedback to improve LLM capabilities (Wu et al., 2023; Grunde-McLaughlin et al., 2023). There is even opportunity for LLM-backed agents to reproduce group- or community-level social insights, with some works examining cultural patterns by interrogating LLM outputs, while others use LLMs to simulate samples of populations via prompting models with “personas” (Aher et al., 2022; Cheng et al., 2023). These advances coupled with demand from communities to scale up decision-making inspired us to additionally evaluate the CLG workflow when LLM prompts are constructed to take the place of human annotators in the human-led process.
However, while LLMs may be able to capture some aspects of social context, they are also susceptible to biases with simulated personas behaving in overtly stereotypical ways (Gupta et al., 2024). To address some of the deficiencies of LLMs, retrieval augmented generation (Lewis et al., 2020b, a) (RAG) has been proposed which combine retrieval from knowledge bases of facts with reasoning capabilities of LLMs and prompt chains, to reduce model reliance on its own knowledge distributions, which may be biased, inaccurate, or out-of-date. While most applications of RAG have been centered around (factual) knowledge retrieval to combat hallucinations (Béchard and Ayala, 2024; Lewis et al., 2020a), recent work has also looked at retrieval over social contexts like personalization (Baek et al., 2024). The increasing recognition of the value of dynamic retrieval of past examples to introduce external factual knowledge in LLM space also nicely parallels ideas in CLG where precedents are retrieved to introduce knowledge around group and community norms and values to ground subjective socially-situated decisions.
Finally, while prompting has also been applied to the task of aligning the models themselves, through approaches like in-context learning (Han, 2023; Hendrycks et al., 2020) which incorporate static few-shot examples as a part of system prompts to guide model behavior, how models apply this information towards generated output can be opaque and unreliable (Arditi et al., 2024). Applying precedents this way would potentially run into similar issues around consistency as constitutional approaches with human-led decision-making. Thus, rather than rely on the models to correctly synthesize decisions by itself through examples, we implement the LLM-prompting version of CLG in a way that follows the human-led version—limiting model decisions to selection of precedents rather than and synthesizing the final decision directly.

### v4

@@ Abstract.
From moderating content within an online community to producing socially-appropriate generative outputs, decision-making tasks—conducted by either humans or AI—often depend on subjective or socially-established criteria. To ensure such decisions are consistent, prevailing processes primarily make use of high-level rules and guidelines to ground decisions, similar to applying “constitutions” in the legal context. However, inconsistencies in specifying and interpreting constitutional grounding can lead to undesirable and even incorrect decisions being made. In this work, we introduce “case law grounding” (CLG)—an approach for grounding subjective decision-making using past decisions, similar to how precedents are used in case law. We present how this grounding approach can be implemented in both human and AI decision-making contexts, introducing both a human-led process and a large language model (LLM) prompting setup. Evaluating with five groups and communities across two decision-making task domains, we find that decisions produced with CLG were significantly more accurately aligned to ground truth in 4 out of 5 groups, achieving a 16.0–23.3 %-points higher accuracy in the human process, and 20.8–32.9 %-points higher with LLMs. We also examined the impact of different configurations with the retrieval window size and binding nature of decisions and find that binding decisions and larger retrieval windows were beneficial. Finally, we discuss the broader implications of using CLG to augment existing constitutional grounding when it comes to aligning human and AI decisions.
@@
journalyear: 2025copyright: ccconference: Collective Intelligence Conference; August 4–6, 2025; San Diego, CA, USAbooktitle: Collective Intelligence Conference (CI 2025), August 4–6, 2025, San Diego, CA, USAdoi: 10.1145/3715928.3737487isbn: 979-8-4007-1489-4/2025/08ccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
Many technological systems today are entangled in social decision-making situations. Online spaces for inter-personal communication, like social media platforms, need to contend with decisions around how to moderate content in a way that maintains a healthy community (Gongane et al., 2022). General purpose AI assistants, like LLM-backed chatbots, also need to implicitly make decisions around how to generate answers that users find appropriate (Ji et al., 2025; Shen et al., 2023; Gabriel, 2020). The scale of these systems means that getting the decisions wrong can have major negative consequences for people (Haimson et al., 2021; Weidinger et al., 2021).
Unlike systems that work with more objective determinations, decisions around social concepts are especially hard because: (1) decision criteria can vary across different groups and tasks; and (2) it can be difficult (or even impossible) to formalize these criteria accurately. Efforts to produce high-quality decisions often involve significant effort, making use of deliberation (Im et al., 2018; Small et al., 2021), juries of community members (Fan and Zhang, 2020), and even expert panels (Klonick, 2020). Scale also creates additional challenges, ultimately leading to the need for automated tools (Jhaver et al., 2019b; Wulczyn et al., 2017) or paid crowd workers (Satariano and Isaac, 2021) to make final decisions, both of which have introduced inconsistencies and errors.
@@
In order to improve the consistency of decisions, processes often first create artifacts to ground the decision-making, the most common of which takes the form of rules and guidelines (which we collectively refer to as constitutions). For example, moderators of communities on social media platforms often create community guidelines that establish what content is allowable, and how violations are to be dealt with (Chandrasekharan et al., 2018; Chandrasekharan and Gilbert, 2019). Similarly, builders and operators of AI assistant services often come up with principles (Huang et al., 2024) that are then used to synthesize training and fine-tuning examples (Bai et al., 2022) or built into system prompts (Lin et al., 2023). However, rules and criteria make use of imprecise language that can leave room for ambiguity (Fiesler et al., 2018; Chen et al., 2021) around edge cases, and varied decision-makers often have divergent interpretations (Gordon et al., 2021). Is there another way we could better support these subjective decisions at scale?
@@
Taking inspiration from case law in legal systems, we introduce case law grounding (CLG), a new approach to grounding decision-making processes by making use of past decisions (precedents) rather than rules (constitutions) as guidance. CLG accomplishes this through 3 steps: (1) case retrieval, where past cases potentially relevant to the current decision are retrieved; (2) precedent selection, where retrieved precedents are further refined for applicability to the current decision; and finally (3) decision synthesis, where the selected precedents and decisions are aggregated and synthesized into the decision for the new case. To address scaling needs, we implemented this grounding approach both as a human-led version where we support human decision-makers (e.g., members of a community or group), and as a LLM-prompting version where an LLM-backed agent is making the decisions.
@@
We evaluated our system on two decision-making tasks: content moderation, where binary “keep”/“remove” decisions were made on Reddit comments (Park et al., 2021), and toxicity rating, where social media posts were rated on a 5-point scale (Kumar et al., 2021). Recruiting participants from 3 Reddit communities ($n=35$) and 2 demographically-determined population groups ($n=60$), we evaluated the human-led version of CLG against a baseline approach that used established constitutions (community guidelines and toxicity annotation guidelines) to support grounding. Using off-the-shelf LLMs, we also evaluated the LLM-prompting version of CLG against a similar baseline where LLMs conducted decision-making using the same established guidelines.
In summary, we make the following contributions:
  • We introduce a general framework for supporting decision-making in subjective and social contexts, case law grounding (CLG), and implement two versions: a human-led process, and an LLM-prompting process.
@@
  • We evaluate CLG across two decision task domains (spanning a total of 5 different group-aligned decision tasks), comparing CLG against a more traditional approach that uses constitutional grounding. We find that:
@@
    • In both the human-led or LLM-prompting process, CLG was able to produce final decisions that more accurately matched observed ground truth compared to decisions grounded with constitutions, with statistically significant improvements observed in 4 out of 5 group decision tasks. (RQ1)
@@
    • For most domains, larger retrieval windows resulted in higher accuracy, however scaling characteristics varied across groups. (RQ2a)
    • Using binding precedents that directly affect the final decision is important—with the 4 groups showing statistically significant improvements only doing so in the binding version. (RQ2b)
  • We discuss our findings in the broader context of grounding subjective, socially-situated decision-making tasks, noting the benefits and drawbacks of each type of grounding, and point to how case law and constitutional grounding can complement each other in supporting group-aligned subjective decision-making.
@@ 2. Related Work
In this section, we discuss how our work relates to prior work by examining the challenges around subjective and social decision-making, designs for and applications of constitutional grounding, and precedents for cases to aid decision-making in prior social and technical systems.
2.1. Challenges in Grounding Subjective and Socially-Established Concepts
Subjective and socially-defined concepts are common in many situations where humans or AI systems need to make decisions, posing unique challenges. For example, online communities on social media platforms often have group-specific shared knowledge in the form of norms or “common sense” criteria around what content (or behavior) is appropriate (Chandrasekharan et al., 2018). These norms often contribute significantly towards how moderation decisions are made (Matias, 2019), but frequently are not explicitly documented (Juneja et al., 2020). This often makes it challenging for those without such context to accurately judge situations even in the presence of written community guidelines (Jhaver et al., 2019a, c). Additionally, concepts in written guidelines can often be under-specified or even conflicting, which can lead to diverging interpretations (Weld et al., 2024; Gordon et al., 2021). Furthermore, socially-established norms and criteria are often fluid, changing over time as the community acquires new members or demographics shift (Lin et al., 2017). These factors all make it hard to create and maintain artifacts for “grounding” these concepts such that they are understood consistently across people and over time.
Beyond online communities, even annotation tasks can involve some degree of subjectivity. For example, instructions or task designers that seem clear to crowdsourcing requesters may actually be confusing to members of a crowd (Wu and Quinn, 2017). Surveys or rating tasks can also involve questions that are by nature subjective (Bentley et al., 2017), ambiguous, or a combination of both (Chen and Zhang, 2023). Efforts in the crowdsourcing realm to build more effective grounding often end up running into conflicting goals—where concepts need to be both unambiguously specified so the annotation can be done while also leaving just the right amount of flexibility for providing the exact subjectivity that we want to capture.
2.2. Rules, Guidelines, and Constitutions
On big social network platforms that host content for millions or billions of people, consistency and transparency around moderation play a paramount role in user trust, making it important to have a clear set of content guidelines and even training materials that ground moderation decision-making (Guardian, 2017; Klonick, 2020). These sets of constitutional rules are costly to develop and often require training of moderators to execute consistently—largely relegating them to bigger organizations where such costs can be justified (ROBERTS, 2019; Satariano and Isaac, 2021) Even in smaller online communities, rules and community guidelines often serve as ways to ground the criteria for what is acceptable content (Fiesler et al., 2018). However, prior work has also observed that in many cases real-world moderation decisions do not always match the expectations of users even given community guidelines (Haimson et al., 2021; Jhaver et al., 2019a, c).
Indeed, not all constitutional grounding needs to take the form of rules and guidelines applied by humans and thus subject to subjective interpretation. Online communities have built constitutions in the form of code to automatically execute decisions, such as with systems like AutoModerator (Jhaver et al., 2019b), ORES (Halfaker and Geiger, 2020), and PolicyKit (Zhang et al., 2020). Constitutional grounding can also form grounding that is internally applied to ground model behavior, such as with efforts around Constitutional AI (Bai et al., 2022; Huang et al., 2024) where humans provide guidelines that ground the behavior of LLMs. However, even in the case of constitutions implemented as code, models can introduce new biases, and the problem of incomplete or conflicting constitutions can still cause decisions that do not align with the intentions of groups or communities.
2.3. The Case for Cases
@@
Cases and case studies have broadly been used as a medium for reasoning around complex scenarios in various fields similar to the socially-constructed judgments we focus on. For example, moral case deliberation (Pierce, 2013; Sush and Najdowski, 2021; Emelin et al., 2021) is often used in medical settings to reflect on ethical situations (Molewijk et al., 2008), and case studies are common in business educational settings too (Farashahi and Tajeddin, 2018). In legal education, fictional cases are frequently used in exercises as a medium for engaging in deliberative reasoning (Gaubatz, 1981; Ringel, 2004). Case-based reasoning (CBR) that focuses on analyzing the circumstances around specific case scenarios also has a long and rich history as a method for thinking about moral philosophy (Paulo, 2015) and in the realm of jurisprudence in the US legal system (Kagan, 2001; Caputo, 2024).
In crowdsourced annotation, cases in the form of examples can augment rubrics to demonstrate how judgments should be made (Wu and Quinn, 2017), act as tests to evaluate whether a concept or criteria was accurately understood (Liu et al., 2016; Hettiachchi et al., 2021; Singla et al., 2014), and even pose as reference anchors to ground the levels on a subjective scale (Chen et al., 2021). More broadly speaking, a data-centric view around subjective concepts posits that data points can also be broadly interpreted as cases (Simons et al., 2020; Jiang and de Marneffe, 2022), with the training of classifiers using human-annotated data being itself a form of grounding (a model’s) decisions through the medium of cases. Though in many datasets, issues around the original data collection often limit their utility as a good model of subjectivity (Arhin et al., 2021; Denton et al., 2021).
@@
Finally, the detail provided by individual cases means that they can act as a medium for deliberation to reconcile differences and work towards shared understanding. We can see examples of this in the legal realm around adjudication (Carpenter, 1917), medical realm for diagnosis (Bertens et al., 2013; Schaekermann et al., 2019), in communities for working out moderation policy (Fan and Zhang, 2020; Pan et al., 2022), and more generally just around challenging tasks involving different lines of reasoning (Chen et al., 2019).